OpenStudy (anonymous):

Find the four arithmetic means between 12 and -13

OpenStudy (anonymous):

between 12 and -13?

OpenStudy (tkhunny):

12 12+a 12+2a 12+3a 12+4a 12+5a = -13 Go!

OpenStudy (anonymous):

the options are A. 6,-1,-3,-8 B.6,2,-3,-8 C. 7,-1,-3,-8 D. 7,2,-3,-8

OpenStudy (whpalmer4):

Solve 12 + 5a = -13. With the value of a (the constant difference between each term), you can find the four arithmetic means between 12 and -13.

OpenStudy (whpalmer4):

They will simply be: 12+a = 12+2a = 12+3a = 12+4a = Just like tkhunny said.